OpenStudy (anonymous):

factor the expression 2x^2-5x+1

OpenStudy (campbell_st):

well you could use x(2x -5) + 1 thats about the sum total of things..

OpenStudy (anonymous):

do u know how to complete the square?

OpenStudy (anonymous):

\[2x ^{2}-5x+1\] \[2(x ^{2}-\frac{ 5 }{ 2 }x ^{2}+\frac{ 1 }{ 2 })\]

OpenStudy (anonymous):

\[2[(x-\frac{ 5 }{ 4 })^{2}-(\frac{ 5 }{ 4 })^{2}+\frac{ 1 }{ 2 }]\]

OpenStudy (anonymous):

\[2(x-\frac{ 5 }{ 4 })^{2}-\frac{ 17 }{ 8 }\]